### v2

# Condense, Don’t Just Prune: Enhancing Efficiency and Performance in MoE Layer Pruning

Abstract
Mixture-of-Experts (MoE) has garnered significant attention for their ability to scale up neural networks while utilizing the same or even fewer active parameters. However, MoE does not relieve the massive memory requirements of networks, which limits their practicality in real-world applications, especially in the era of large language models (LLMs). While recent work explores the possibility of removing entire layers of MoE to reduce memory, the performance degradation is still notable. In this paper, we propose ConDense-MoE (CD-MoE) that, instead of dropping the entire MoE layer, condenses the large, sparse MoE layer into a smaller, denser layer with only a few experts activated for all tokens, while maintaining hardware friendliness. Our approach is specifically designed for fine-grained MoE with shared experts, where Feed-Forward Networks are split into many small experts, with certain experts isolated to serve as shared experts that are always activated, such as DeepSeekMoE and QwenMoE. We demonstrate the effectiveness of our method. Specifically, for the DeepSeekMoE-16B model, our approach maintains 90% of the average accuracy while reducing memory usage by 27.5% and increasing inference speed to 1.26 times. Moreover, we show that by applying lightweight expert fine-tuning—only to the condensed layers—and using 5 hours on a single 80G A100 GPU, we can successfully recover 98% of the original performance. Our code is available at https://github.com/duterscmy/CD-MoE/tree/main.
Condense, Don’t Just Prune:
Enhancing Efficiency and Performance in MoE Layer Pruning
Mingyu Cao1, Gen Li2, Jie Ji2, Jiaqi Zhang3, Xiaolong Ma2, Shiwei Liu4, Lu Yin5thanks: Correspondence to Lu Yin: l.yin@surrey.ac.uk 1Shopee Pte. Ltd, China  2Clemson University, USA  3Meituan Inc, China 4University of Oxford, United Kingdom  5University of Surrey, United Kingdom
Figure 1: Left: The structure of the Deepseek MoE layer. $w_{i}^{\prime}$ represents the weights after normalization. Right: The structure of the ConDense-MoE layer, where only the most important top-k experts are retained. $\bar{w_{i}}$ represents the fixed weights that are pre-computed during condensing using the average weight of all calibration tokens.
1 Introduction
Large Language Models (LLMs) continue to grow in both size and capability (Brown, 2020; Touvron et al., 2023b, a), fueling demand for architectures that can scale with minimal increases in computational cost. Mixture of Experts (MoE) architectures have emerged as a promising solution to this challenge (Shazeer et al., 2017; Jiang et al., 2024a; Team et al., 2024). Unlike standard dense networks, MoEs selectively activate only the most relevant subset of parameters (referred to as “experts”) for a given input, enabling substantial growth in model capacity without a proportional escalation in compute overhead. Recent advances in fine-grained expert segmentation (Dai et al., 2024; Team, 2024) push this concept further by splitting feed-forward layers into many small experts, while designating a small set of shared experts that remain active across all tokens. Despite the clear benefits, MoE architectures still face major deployment barriers due to high memory costs. Storing a large number of experts, even if most remain dormant per token, can be prohibitively expensive in real-world systems. This reality highlights the urgent need for approaches that preserve the performance benefits of MoEs while lowering their memory footprint.
In this paper, we present a new framework called ConDense-MoE (CD-MoE), a novel framework that significantly enhances the efficacy specifically for DeepSeek-style shared-expert-type MoE (Dai et al., 2024; Team, 2024). Unlike prior MoE compression approaches that often remove entire layers or retain the routing mechanism with fewer experts (He et al., 2024), our method eliminates the routing process at selected layers and condenses them into dense layers. Specifically, we devise a greedy strategy to identify which layers—and which experts within them—can be condensed with minimal impact on model accuracy, leading to significant reductions in memory usage and faster inference.
CD-MoE stems from a surprising yet critical insight: in shared-expert–type MoEs, simply removing the routing mechanism while retaining only the shared experts at a given layer generally incurs minor output difference. Building on this, CD-MoE preserves these essential shared experts and a small subset of routing experts, thereby condensing originally sparse layers into compact, dense analogs. As these preserved experts account for only a small fraction of overall parameters, we substantially cut memory requirements without sacrificing much performance. Our experiments demonstrate that on DeepSeekMoE-16B (Dai et al., 2024), CD-MoE reduces memory by 27.5% and maintains nearly 90% of the original accuracy on a suite of downstream tasks. Moreover, our method uncovers additional experts that are beneficial for fine-tuning, enabling up to 98% recovery of the original performance after a lightweight supervised fine-tuning that updates only the condensed layers. In summary, our contributions are:
  • A novel MoE condensation approach that selectively removes less critical experts and condenses large sparse MoE layers into small, dense structures.
  • An efficient selection algorithm that identifies which layers to condense and which experts to preserve, achieving a favorable trade-off between memory savings and accuracy.
  • Empirical validation on DeepSeekMoE-16B, demonstrating a 27.5% memory reduction and a 1.26$\times$ inference speed, while retaining up to 90% of the zero-shot accuracy. Moreover, lightweight fine-tuning can effectively reclaim 98% of the original performance using only a single A100 GPU in a few hours.
Input: Calibration data input $X$, routing expert set $E_{\text{route}}=\{E_{1},\ldots,E_{n}\}$, number of experts to select $K$
Output: Condense expert set $E_{\text{condense}}=\{E_{\text{shared}}\}$
Initialize $E_{\text{condense}}\leftarrow\{E_{\text{shared}}\}$,  $O_{\text{ref}}\leftarrow\text{Layer}_{\text{route}}(X)$trianglerightCompute reference output with all routing experts
for $k\leftarrow 1$ to $K$ do
       for each expert $E_{i}\in E_{\text{route}}$ do
             $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{i}\}$trianglerightTemporarily add expert to condense set
             $O_{\text{tmp}}\leftarrow\text{Layer}_{\text{condense}}(X)$trianglerightCompute output with current condense experts
             $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$trianglerightCompute JS divergence as loss for $E_{i}$
             $E_{\text{condense}}\leftarrow E_{\text{condense}}\setminus\{E_{i}\}$trianglerightRemove expert after computing loss
      $E_{\text{best}}\leftarrow\text{argmin}_{E_{i}\in E_{\text{route}}}\,loss_{i}$trianglerightSelect the expert with minimum loss
       $E_{\text{condense}}\leftarrow E_{\text{condense}}\cup\{E_{\text{best}}\}$trianglerightAdd best expert to condense set
       $E_{\text{route}}\leftarrow E_{\text{route}}\setminus\{E_{\text{best}}\}$trianglerightRemove selected expert from routing set
Algorithm 1 Condense Experts Selection
2 Related Work
Mixture-of-Experts (MoE) architectures have become a powerful framework for scaling neural networks without linearly increasing computational burden (Shazeer et al., 2017; Fedus et al., 2022; Jiang et al., 2024a; Dai et al., 2024; Team, 2024). Early methods such as Switch Transformer (Fedus et al., 2022) demonstrated that sparse activation can substantially boost model capacity while maintaining tractable compute. Further work has refined routing policies (Jiang et al., 2024b) and introduced a fine-grained segmentation of feed-forward layers into many small experts, often augmented with “shared” experts that remain active for every token (Dai et al., 2024; Team, 2024). These advances collectively enhance training stability and generalization, yet the growing pool of experts poses serious deployment challenges, chiefly due to large memory demands.
A range of pruning and compression strategies have emerged to address these overheads. Some works prune experts by identifying those that are minimally activated or less essential for downstream tasks (Chi et al., 2022; Sun et al., 2024; Muzio et al., 2024), with solutions ranging from exhaustive searches over all expert subsets (Lu et al., 2024) to filtering experts by their activation frequency. While these methods typically preserve routing, more aggressive approaches prune entire MoE layers, achieving considerable memory reductions but at the risk of pronounced accuracy losses (He et al., 2024). In contrast, our technique selectively eliminates the routing mechanism in particular layers and retains only a small subset of experts (including the shared experts), thus reducing memory demands while sustaining most of the model’s representational power.
3 ConDense-MoE (CD-MoE)
In this section, we present a comprehensive explanation of the CD-MoE process, which follows a sequential workflow comprising several critical steps. First, we delve into Expert Selection and Condensing (Section 3.2), where we detail the methods used to identify and condense the most essential experts from all available experts within a layer. Second, we explore Layer Selection and Condensing (Section 3.3), focusing on selecting the most suitable layers for condensation to preserve as much of the LLM’s inherent capabilities as possible. Last but not least, we highlight that only performing lightweight fine-tuning (Section 4.3) on condensed layers, can almost recover the original performance of MoE.
Input: Data input $X$, routing layer set $L_{\text{route}}=\{L_{1},\ldots,L_{n}\}$, number of layers to select $K$
Output: Condense layer set $L_{\text{condense}}=\{\}$
Initialize $L_{\text{condense}}\leftarrow\{\}$,  $O_{\text{ref}}\leftarrow\text{Model}(X)$trianglerightCompute reference output with all routing layers
for $k\leftarrow 1$ to $K$ do
       for each layer $L_{i}\in L_{\text{route}}$ do
             $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{i}\}$trianglerightTemporarily add layer to condense set
             $O_{\text{tmp}}\leftarrow\text{Model}(X)$trianglerightCompute output with current condense layers
             $loss_{i}\leftarrow\text{JS}(O_{\text{ref}},O_{\text{tmp}})$trianglerightCompute JS divergence as loss for $L_{i}$
             $L_{\text{condense}}\leftarrow L_{\text{condense}}\setminus\{L_{i}\}$trianglerightRemove layer after computing loss
      $L_{\text{best}}\leftarrow\text{argmin}_{L_{i}\in L_{\text{route}}}\,loss_{i}$trianglerightSelect the layer with minimum loss
       $L_{\text{condense}}\leftarrow L_{\text{condense}}\cup\{L_{\text{best}}\}$trianglerightAdd best layer to condense set
       $L_{\text{route}}\leftarrow L_{\text{route}}\setminus\{L_{\text{best}}\}$trianglerightRemove selected layer from routing set
Algorithm 2 Condense Layers Selection
3.1 Preliminaries of MoE
Here we define the basic structure of an MoE layer. All experts in an MoE layer are represented as $\{E_{1},E_{2},\dots,E_{n}\}$. For most fine-grained MoE models, in addition to routing experts, the shared experts $E_{s}$ is used. Any token will activate the shared experts and several routing experts, as shown in Figure 1 left. For the input data $X$, the output is computed as follows:
$\mathbf{h}_{t}=\sum_{i=1}^{N}\left(g_{i,t}\,E_{i}\left(\mathbf{x}_{t}\right)\right)+E_{s}\left(\mathbf{x}_{t}\right)+\mathbf{x}_{t},$ (1)
where
$g_{i,t}=\begin{cases}s_{i,t},&s_{i,t}\in\text{TopK}\left(\{s_{j,t}\mid 1\leq j\leq N\},K\right),\\ 0,&\text{otherwise},\end{cases}$
$s_{i,t}=\text{Softmax}_{i}\left(\mathbf{x}_{t}^{T}\mathbf{e}_{i}\right).$
Here, $N$ denotes the total number of route experts, $E_{i}(\cdot)$ represents the $i$-th expert, $E_{s}$ is the shared expert, $g_{i,t}$ is the gate value for the $i$-th expert, $s_{i,t}$ denotes the token-to-expert similarity, $\text{TopK}(\cdot,K)$ represents the set consisting of the $K$ highest similarity scores among those computed for the $t$-th token and all $N$ experts, and $\mathbf{e}_{i}$ is the centroid of the $i$-th expert.
3.2 Expert Selection and Condensing
Our CD-MoE framework removes the standard routing mechanism during inference. For each expert $E_{i}$, we use C4 calibration data to inference and compute the mean of $g_{i,t}$ when that expert is activated, which is then used as the fixed gate value $g_{i}$ for that expert during the inference phase:
$g_{i}=\frac{1}{|\{t\mid g_{i,t}\neq 0\}|}\sum_{t\mid g_{i,t}\neq 0}g_{i,t}$ (2)
For the condense layer, we only retain $K$ routing experts including the shared experts, which is consistent with the number of active experts during the training phase. Additionally, benefiting from sufficient training during the training stage, the shared expert is retained, as shown in Figure 1 right. After condensation, the MoE layer is represented as:
$\mathbf{h}_{t}=\sum_{i=1}^{K}\left(g_{i,t}\,E_{i}\left(\mathbf{x}_{t}\right)\right)+E_{s}\left(\mathbf{x}_{t}\right)+\mathbf{x}_{t}$ (3)
Next, we describe how to select the most crucial experts. Lu et al. (2024) explore all possible expert combinations and choose the combination of experts that makes the layer outputs before and after pruning as close as possible, while pruning the rest. However, for fine-grained MoE models like DeepSeekMoE-16B (Dai et al., 2024) and Qwen1.5-MoE-A2.7B (Team, 2024), where the number of experts often exceeds 60, testing all permutations of experts requires significant computational resources and time. Therefore, we propose using a greedy search approach Algorithm 1 to select the most critical experts.
In which, ${JS}(.)$ is the Jensen-Shannon divergence used to measure the difference between two outputs. Zhang et al. (2024) shown that in model pruning, this metric is better than angular distance and Euclidean distance. We also demonstrated its superiority over other metrics such as Kullback–Leibler (KL) divergence and perplexity in Table 2. Therefore, we choose ${JS}(.)$ divergence as our selection metric. The mathematical expression of ${JS}(.)$ is as follows:
$JS(u,v)=\frac{1}{2}\left(KL\left(u\,,\,\frac{1}{2}(u+v)\right)\right.\\ \left.+KL\left(v\,,\,\frac{1}{2}(u+v)\right)\right)$ (4)
Here, ${KL}(.)$ represents the Kullback–Leibler divergence, defined for discrete random variables as:
$KL(u,v)=\sum_{j}u_{j}\log\left(\frac{u_{j}}{v_{j}}\right)$ (5)
where $u$ and $v$ are vectors that represent discrete probability distributions, which, in our method, can be the dense layer output and the condensed layer output.
3.3 Layer Selection and Condensing
Clark et al. (2019b) have confirmed that different layers in language models exhibit distinct functionalities. While our objective is to retain the most impactful experts during the condensing process of each layer, it is inevitable that the outputs will deviate from their original values due to reduced capacity. Indiscriminately condensing all layers can thus lead to significant degradation in model performance and the loss of essential features. Therefore, akin to our approach in expert selection, we prioritize condensing those layers that exert minimal impact on model output changes post-pruning.
Figure 2: Fluctuations in the $JS$ divergence between the the outputs of the condensed model and the original dense model across different layers.
To systematically quantify the impact of condensing individual layers, we conducted preliminary experiments assessing the output changes using the Jensen-Shannon ($JS$) divergence between the pruned and unpruned model outputs as our evaluation metric. As illustrated in Figure 2, condensing layer 0 results in a significant shift in the output distribution, indicating its critical role in the model’s performance. While the middle layers exhibit some fluctuation, condensing the last three layers also causes noticeable deviations, suggesting that these layers are essential for maintaining output fidelity. These observations indicate that condensing different layers introduces varying degrees of change in the $JS$ metric, reflecting their varying importance to the overall model performance.
Table 1: Comparison with the dense models, unpruned MOE model and the baseline methods, the underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#$\#L$’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model. CD-MoE-S refers to retaining only the Shared expert after condensing, while CD-MoE-SR means selecting additional 6 Routed experts and keeping as dense.
method #l act. mem. speedup arc-c boolq hella mmlu obqa piqa rte wino avg.
gpt-neo-2.7b - 2.7B - - 29.8 61.7 55.2 25.0 34.6 72.9 53.4 57.8 48.8
opt-2.7b - 2.7B - - 31.2 59.9 60.6 25.4 35.2 74.7 54.2 60.5 50.2
bloom-3b - 3B - - 30.4 61.4 54.5 25.9 32.4 70.7 56.0 58.6 48.7
openllama-3b - 3B - - 36.1 67.1 64.4 23.9 38.6 75.1 54.2 62.3 52.7
DeepseekMoE-16b - 2.8B 100% 1.0x 48.3 72.7 77.4 38.3 44.2 78.7 63.2 70.1 61.6
w/o sft
LayerTrim 8 2.4B 72.2% 1.34x 35.3 65.4 57.5 29.8 32.6 62.9 48.0 63.6 49.4
BlockTrim 8 2.3B 71.3% 1.42x 36.2 62.2 57.2 29.0 34.6 70.2 63.2 64.7 52.2
CD-MoE-S 8 2.4B 73.1% 1.34x 37.8 70.6 65.9 28.8 38.6 75.0 58.8 65.0 55.1
CD-MoE-SR 9 2.8B 72.5% 1.26x 37.9 72.3 64.4 27.1 37.6 75.9 61.7 67.0 55.5
w/ lightweight sft
CD-MoE-S + SFT 8 2.4B 73.1% 1.34x 41.3 72.1 65.9 39.7 40.0 76.2 67.9 69.7 59.1
CD-MoE-SR + SFT 9 2.8B 72.5% 1.26x 42.6 79.2 66.9 38.6 38.6 75.7 66.8 75.2 60.5
Motivated by these findings, we adopt a greedy search strategy analogous to that used in expert selection to determine the optimal layers for condensing. Specifically, we compute the JS divergence before and after condensing based on the output of the final layer, iteratively adding layers to the $Layer_{condense}$ set that can minimize the distance between the model’s final layer output and the output before condensing, as detailed in Algorithm 2. For layers that are not condensed, the standard token routing mechanisms are employed during forward propagation to maintain their full expressive power. For the layers being condensed, the routing mechanism is removed, and the MOE is compressed into a dense structure. This selective condensing approach allows us to effectively reduce computational complexity and memory usage while minimizing degradation in model performance.
4 Experiments
4.1 Experiment Setup
We adopt the opensource model DeepSeekMoE-16B and Qwen1.5-MoE-A2.7B (Team, 2024). Besides, in alignment with the methodologies outlined by Sun et al. (2024) and Lu et al. (2024), we randomly sample 100 examples from the C4 dataset to serve as calibration data during the condensing phase of our model.
Evaluation: To provide a comprehensive assessment of our model’s performance, we follow the protocol established by He et al. (2024). We report zero-shot accuracies on eight diverse tasks selected from the EleutherAI Language Model Evaluation Harness (Gao et al., 2024), which includes the 7 commonsense reasoning benchmark datasets: ARC-Challenge (Clark et al., 2018), BoolQ (Clark et al., 2019a), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2019), RTE (Dagan et al., 2005), WinoGrande (Sakaguchi et al., 2019), and a more challenging dataset MMLU (Hendrycks et al., 2021).
Baseline: For comparative analysis, we primarily benchmark our approach against the Block Trimming and Layer Trimming methods proposed by He et al. (2024). These methods serve as strong baselines for pruning MoE models and are directly relevant to our study. Both pruning methods prune all experts in the layer. They adopt Block Influence for pruning as introduced in Men et al. (2024).
We also compared the results with dense models that have similar activations, such as GPT-Neo-2.7B (Gao et al., 2020), OPT-2.7B (Zhang et al., 2022), BLOOM-3B (Press et al., 2022), and OpenLLaMA-3B (Touvron et al., 2023a).
Figure 3: CD-MoE against baselines on zero-shot tasks w/o fine-tuning. Left: Average accuracy with varying Memory Ratio against the original model. Right: Average accuracy with varying SpeedUp against the original model. The Gray dotted line is the original model result. CD-MoE-S represents the shared experts and no routing experts, and CD-MoE-SR represents shared with routing experts. Baseline indicated performance of the dense model.
4.2 Zero-Shot Evaluation
To assess the effectiveness of our condensation approach in a purely zero-shot setting, we evaluate CD-MoE-S and CD-MoE-SR on eight downstream tasks without any additional fine-tuning. As shown in Table 1, we report each model’s performance alongside its memory reduction ratio. Because pruning different layers can lead to varying memory footprints, we normalize the number of pruned layers so that all methods operate under comparable memory budgets.
Specifically, ❶ Comparison to Baselines: Under similar memory constraints, both CD-MoE-S and CD-MoE-SR consistently outperform the layer-pruning baselines—BlockTrim and LayerTrim—highlighting the importance of preserving the most informative MoE experts. Moreover, Figure 3 plots the average accuracy against increasing pruning rates and memory savings, illustrating that the performance gap between CD-MoE and conventional pruning techniques widens as pruning becomes more aggressive. ❷ Comparison to Small Dense LLMs: Notably, even without fine-tuning, CD-MoE-S and CD-MoE-SR often surpass smaller dense LLMs of comparable size (e.g., BLOOM-3B, OpenLLaMA-3B). This suggests that training a dense model from scratch at a reduced scale does not match the performance of selectively condensed MoE architectures that retain high-value parameters. ❸ CD-MoE-S vs. CD-MoE-SR: In zero-shot mode, CD-MoE-S exhibits comparable or slightly superior performance to CD-MoE-SR. However, as detailed in Section 4.3, the routed experts in CD-MoE-SR confer a substantial advantage during fine-tuning: they enable nearly 98% of the original performance to be recovered in just a few hours on a single 80GB A100 GPU. The results for Qwen1.5-MoE-A2.7B is reported in Appendix A.4.
Overall, these results confirm that our condensation strategy preserves core model capacities more effectively than layer-level pruning alone. By selectively retaining expert parameters, CD-MoE consistently delivers strong zero-shot accuracy while substantially reducing memory and computational overhead.
4.3 Lightweight Fine-tuning Evaluation
Figure 4: CD-MoE with lightweight fine-tuning. Left: SFT results on CD-MoE-S with increasing number of condensed layers. Right: SFT results on CD-MoE-SR with increasing number of condensed layers. Baseline indicated performance of the dense model.
We further experimented with expert fine-tuning and improving performance on specific tasks.
Since the condensed model inevitably loses part of its language modeling capability, further fine-tuning can restore the model’s abilities (Shi et al., 2023; Sanh et al., 2020; Frantar and Alistarh, 2023). Our goal is to demonstrate that fine-tuning can restore the capabilities of the condensed layers to their pre-condensation state, so we propose a lightweight fine-tuning approach. Specifically, only the parameters in the condensed layers are updated to conduct efficient fine-tuning. For example, with CD-MoE-SR, only 3.8% of the parameters update gradients, and it takes just 2 hours on a single A100 GPU to perform SFT on the MMLU training set. For SFT, following the approach of Li et al. (2024), we fine-tuned using commonsense170k and MMLU training set respectively.
Table 1 (bottom panel) summarizes the performance gains attributable to lightweight SFT. CD-MoE-S improves from an average score of 55.1 to 59.1, while CD-MoE-SR jumps from 55.5 to 60.5—nearly matching the original uncondensed performance of 61.6. Notably, ❶ Commonsense Reasoning: Task like BoolQ shows marked improvements. The CD-MoE-SR + SFT lifts performance by up to 7 points in compared to its counterpart before fine-tuning. ❷ Complex Knowledge Tasks: On MMLU, we observe significant gains (e.g., from 27.1 to 38.6 with CD-MoE-SR + SFT), indicating that even after pruning, the model retains foundational knowledge that can be effectively recovered through selective fine-tuning. ❸ Efficiency: Despite the notable performance boost, the computational overhead remains low because only a fraction of parameters ($\sim 3\%$) is updated. This design ensures training remains practical for resource-constrained settings, aligning with real-world demands.
Figure 4 shows how the average accuracy changes after SFT as more layers are condensed. Compared to CD-MoE-S, CD-MoE-SR delivers consistently higher accuracy with lower memory usage, suggesting that selecting additional routing experts via greedy search preserves more of the model’s foundational knowledge. Moreover, CD-MoE-SR can restore and even surpass the original model’s performance through lightweight SFT, reducing memory usage by up to 20%. These findings indicate that our condensation strategy retains recoverable knowledge despite significant pruning. With targeted lightweight SFT, the condensed experts regain most of their original performance, and in certain tasks, CD-MoE-SR + SFT even exceeds the unpruned baseline.
4.4 Experiment Analysis
Expert selection algorithms. We evaluate several strategies for selecting the most critical experts at each layer. Our primary baselines rely on easily computable statistical properties of the expert parameters:
  • Random: Randomly select $K$ experts.
  • $L_{1}$: Calculate the sum of the $L_{1}$ norms of each expert’s parameters, sort them, and select the $K$ experts with the lowest $L_{1}$ norms.
  • PL_Alpha_Hill: Following heavy-tailed self-regularization Martin and Mahoney (2019, 2020, 2021), we measure each layer’s spectral density via the Hill estimator Hill (1975); Xiao et al. (2023):
[…]
We also experimented with selecting experts having the highest $L_{1}$-norms, but observed substantially worse performance, potentially because lower norms often indicate better-converged parameters. In Table 2, we compare these methods to our Greedy Search, which iteratively selects experts that minimize the deviation between the post-condensation and pre-condensation outputs of the layer. Notably, none of the statistical baselines (Random, L1, or PL_Alpha_Hill) outperform Greedy Search. We hypothesize that the latter’s layer-level output fidelity objective directly retains the most impactful experts.
Table 2: Comparison of expert selection methods. $L_{n}$ denotes condensing $n$ layers. Each entry is the average accuracy (%) on eight downstream tasks.
Methods $L_{6}$ $L_{9}$ $L_{12}$ $L_{15}$
Random 54.6 52.7 50.1 46.8
PL_Alpha_Hill 56.8 52.5 50.3 47.3
L1 56.5 53.2 50.4 47.5
Greedy Search 56.5 55.5 52.8 49.0
We also measure the time cost of our Greedy Search on the DeepSeekMoE-16B model and compare it with the exhaustive traversal in Lu et al. (2024), which enumerates all expert combinations. Each experiment uses 100 C4 calibration samples with maximum sequence length 512 and batch size 256, running on an 80G A100 GPU. As shown in Table 3, the exhaustive approach requires an intractably large number of inferences and is therefore infeasible. By contrast, Greedy Search requires only a few hundred inferences per layer, striking a favorable balance between accuracy preservation and runtime efficiency.
Table 3: Time cost comparison for expert and layer selection. ‘-’ indicates that the method does not require the step, ‘*’ denotes estimates.
Method Expert Layer
Inferences (Counts) Time (Hours) Time (Hours)
Lu et al. (2024) 75,041,808 2500* -
CD-MoE-S - - 0.12
CD-MoE-SR 369 0.13 0.15
Layer selection algorithms. We compare our proposed Greedy Search (Layer) method against two baselines that rank layers by measuring the divergence between the layer outputs before and after condensation:
  • Layer Rank: For each layer, we compute the $JS$ distance between its output before and after condensation. We then sort the layers and choose the one with the smallest divergence, indicating minimal impact on the layer-level output.
  • Global Layer Rank: For each layer, we apply condensation and evaluate the $JS$ distance between the model’s final outputs before and after condensation. Layers are then sorted accordingly, and the layer that most preserves the final output distribution is preferred.
Table 4 shows a comparison of these methods in the context of CD-MoE-SR. Notably, our Greedy Search consistently surpasses the baselines, underscoring its effectiveness in identifying layers that minimally disrupt the overall model output. We also measure the time needed to condense 15 of the 27 MoE layers using the same setup as in expert selection (see Table 3). This operation requires only $0.15$ hours, further validating the computational efficiency of the CD-MoE approach.
Table 4: Comparison of layer selection methods. $L_{n}$ indicates the condensation of $n$ layers. Entries are average accuracies (%) on eight test datasets.
Methods $L_{6}$ $L_{9}$ $L_{12}$ $L_{15}$
Layer Rank 45.5 44.2 42.8 42.2
Global Layer Rank 55.7 53.1 48.8 45.9
Greedy Search 56.5 55.5 52.8 49.0
Searching metrics.
To quantify how well each condensed model approximates the original model’s output, we also compare several metrics during the Greedy Search process. Following Zhang et al. (2024), we adopt $JS$ divergence as our primary measure, as it has been shown to be more sensitive to output changes than angular distance or Euclidean distance. We compare $JS$ against two alternative metrics:
  • Perplexity (PPL): For each calibration sample, we compute the change in PPL and select layers that minimize the average difference in PPL between the original and condensed models.
  • KL Divergence: We measure the shift between the two output distributions using $KL$ divergence and select layers with minimal $KL$ distance.
Table 5 presents the results. While $PPL$ achieves a marginally better score for $L_{6}$, $JS$ divergence excels as more layers are condensed ($L_{9},L_{12},L_{15}$). These findings substantiate the choice of $JS$ divergence as a robust metric for maintaining the model’s overall quality under heavier compression.
Table 5: Performance of different metrics for layer selection. $L_{n}$ indicates the number of condensed layers. The values shown are average accuracies (%) on eight test datasets.
Metrics $L_{6}$ $L_{9}$ $L_{12}$ $L_{15}$
PPL 58.7 55.4 50.9 47.2
KL divergence 57.1 54.7 52.2 48.5
JS distance 56.5 55.5 52.8 49.0
5 Conclusion
We have presented CD-MoE, a novel framework for compressing large-scale Mixture-of-Experts (MoE) models by selectively condensing them into denser layers. Our approach removes the token routing mechanism and uses a greedy search algorithm to prune the majority of less significant experts, allowing all tokens to activate only a handful of highly impactful experts. Extensive experiments on DeepSeekMoE-16B and other fine-grained MoE architectures show that CD-MoE can retain up to 90% of the original model’s accuracy while enjoy 27.5% memory reduction and a 1.26$\times$ inference speed. Moreover, we demonstrate that lightweight expert fine-tuning, focused solely on the condensed layers, can further reclaim the original model’s performance within a few hours on a single GPU. These results highlight the practical benefits of CD-MoE for resource-constrained scenarios, where memory and computational efficiency are critical. Future directions include exploring the synergy of CD-MoE with quantization and knowledge distillation, and extending the method to more diverse MoE structures lacking shared experts.
6 Limitations
While the proposed ConDense-MoE (CD-MoE) method shows encouraging results, it is primarily designed for fine-grained MoE models with shared experts. In more traditional MoE architectures that do not use shared experts, the condensation process may lead to some performance changes due to fewer retained experts.
7 Potential Risk
CD-MoE may raise some ethical considerations related to bias and fairness. Due to the expert selection process, there is a possibility that certain experts or layers may be underrepresented, which could impact the model’s performance or outcomes, especially in cases where diversity and inclusivity are important. While this may not be a significant issue in many scenarios, it is something to keep in mind when applying the method to more sensitive or varied tasks.
Appendix A Appendix
A.1 Fluctuations in Metrics across Condensing Different Layers
In Figure 1, we show that the $\mathrm{JS}$ divergence between the model’s final output and the output before condensing varies substantially across layers. Figure 5 further demonstrates how $\mathrm{KL}$ divergence and perplexity (PPL) change when pruning different layers. Notably, condensing layer 0 continues to produce a significant shift in the model’s outputs, while the middle layers exhibit fluctuations, and the final layers show an increase. Across all layers, the trends in $\mathrm{JS}$, $\mathrm{KL}$, and PPL remain similar.
Figure 5: Left: fluctuations in the $KL$ divergence. Right: fluctuations in the perplexity.
A.2 Implementation Details
We utilize Hugging Face and PyTorch for the implementation of our work  [^1] Our repository is built on top of Transformers: https://github.com/huggingface/transformers. All inference and fine-tuning operations are executed using bf16 (Brain Floating Point 16) precision on NVIDIA A100 GPU equipped with 80GB of memory. During the fine-tuning phase, we employ an initial maximum learning rate of $1\times 10^{-4}$, implement a warm-up ratio of 10% to gradually ramp up the learning rate, and utilize cosine annealing as the learning rate scheduler to ensure smooth convergence and prevent potential overfitting. Due to the large computational requirements and our limited resources, all experiments were conducted in a single run.
The expert configurations of DeepSeekMoE-16B and Qwen1.5-MoE-A2.7B is represented at Table 6.
Table 6: Expert Configurations in MoE Models
Model DeepSeekMoE-16B Qwen1.5-MoE-A2.7B
Parameters 16,375,728,128 14,315,784,192
Activated Parameters 2,828,650,496 2,689,177,536
Number of Experts 66 64
Number of Activated Experts 8 8
Ratio of Activated Experts 12.1% 12.5%
Number of Shared Experts 2 4
Ratio of Shared Experts in Activated Experts 25% 50%
A.3 Calibration Data Analysis
C4 dataset commonly serves as the calibration benchmark. In our study, we aim to ensure that the performance of the condensed model aligns closely with specific task requirements. To this end, we experiment by utilizing questions extracted from multiple test sets as calibration data, thereby tailoring the language modeling capabilities of the condensed model to the tasks at hand. Specifically, we sampled 20 questions from each of eight distinct test sets, resulting in a total of 160 questions.
Figure 6: We compare the model performance between using C4 and downstream task data as calibration data across different layer pruning indexes, under CD-MoE-SR setting on DeepSeek-16B.
As demonstrated in Figure 6, it is noteworthy that across varying numbers of condensed layers, the C4 data consistently achieves superior performance compared to the downstream tasks data. This suggests that the diversity and comprehensiveness of the C4 dataset provide a more robust foundation for calibrating the condensed model. Consequently, we adopt C4 as the calibration data in all experiments. This finding is consistent with the work by Bandari et al. (2024), where they perform unstructured pruning of LLM parameters, while we focus on layer condensation.
A.4 Zero-shot Evaluation on Qwen1.5-MoE-A2.7B
Table 7 presents the results of CD-MoE on the Qwen1.5-MoE-A2.7B model. Under similar memory cost, the advantages of CD-MoE over baseline methods are more pronounced, and with more than 20% memory reduction, it retains about 96% of the original model performance. We attribute this to the fact that Qwen1.5-MoE-A2.7B has a larger proportion of shared experts, as shown in Table 6. Our method preserves these experts, allowing the condensed model to retain more knowledge learned during the pre-training phase.
Table 7: Comparison with the dense models, unpruned MOE model and the baseline methods on Qwen1.5-MoE-A2.7B, the underlined variables are tried to be kept consistent for fair comparison, and the bolds represent the best results. ‘#$\#L$’ represents the number of pruned layers; ‘ACT.’ represents the number of activated parameters; and ‘MEM.’ is the remaining memory ratio compared to the original model. CD-MoE-S refers to retaining only the Shared expert after condensing, while CD-MoE-SR means selecting additional 4 Routed experts and keeping as dense.
method #l act. mem. speedup arc-c boolq hella mmlu obqa piqa rte wino avg.
gpt-neo-2.7b - 2.7B - - 29.8 61.7 55.2 25.0 34.6 72.9 53.4 57.8 48.8
opt-2.7b - 2.7B - - 31.2 59.9 60.6 25.4 35.2 74.7 54.2 60.5 50.2
bloom-3b - 3B - - 30.4 61.4 54.5 25.9 32.4 70.7 56.0 58.6 48.7
openllama-3b - 3B - - 36.1 67.1 64.4 23.9 38.6 75.1 54.2 62.3 52.7
Qwen1.5-MoE-A2.7B 2.7B 100% 1.0x 45.0 79.5 77.3 61.2 43.6 80.3 67.9 69.3 65.5
w/o sft
LayerTrim 6 2.3B 78.3% 1.23x 34.4 62.6 63.1 45.4 32.8 73.6 54.5 62.6 53.6
BlockTrim 6 2.2B 76.1% 1.36x 31.3 62.2 57.4 38.4 32.0 70.7 52.7 58.2 50.4
CD-MoE-S 6 2.5B 78.3% 1.22x 41.1 75.8 72.9 56.0 41.0 78.9 67.2 68.1 62.6
CD-MoE-SR 6 2.7B 79.7% 1.19x 40.2 77.2 72.1 54.9 39.4 78.0 71.1 66.4 62.4